# \$*\$ \$*\$ footnotetext: Corresponding Author

### v2

@@ 1 Introduction
[…]
@@
[…]
@@ 3.1 Overall Framework
$\mathcal{L}_{\text{distill}}=-\sum\log\text{sim}(\mathcal{S}(\textbf{x}_{t}),\text{\tt{SigLIP}}(\textbf{X})),$ (2)
@@ 3.2 Training Recipe
Stage-1 Projector 66M WebVid [8] Pandas [22] OmniCorpus [56]
@@
Stage-2 Full Model (w/o VAE) 9M HQ Und. OpenVid-1M [75] VIST [44]
16M HQ Gen. 1.5M Internal Data CoMM [23]
@@
Stage-2. Subsequently, we tune the full model using 9M high-quality multimodal understanding instruction data and 16M high-quality visual generation data filtered from the 66M image-text pairs.
@@ 4.1 Experimental Setup
Datasets. The curated approximately 66M image-text pairs consist of images with a resolution of at least 512 pixels in width and height. The images are filtered from CC12M [13], COYO [12], LAION-Aesthetic-12M∗*∗ ∗*∗ * https://huggingface.co/datasets/dclure/laion-aesthetics-12m-umap and AI synthetic data. The images are recaptioned by LMMs except for the synthetic data. The 9M high-quality multimodal understanding instruction data is curated from Densefusion-1M [58], and LLaVA-OneVision [52].
Implementation Details. The semantic layers $\mathcal{S}(\cdot)$ are pre-distilled from SigLIP-so400m-patch14-384∗*∗ ∗*∗ * https://huggingface.co/google/siglip-so400m-patch14-384 over 200K iterations, using a batch size of 512 and a cosine-scheduled learning rate of 2e-5. During distillation, Eq. 1 is applied to the visual latents with only a probability of 0.3 in the last 20K iterations. The input image resolution of 3D causal VAE encoder with $2\times 2$ patch embedding layer is set as $432\times 432$ to get $729=27\times 27$ visual latents, which matches the ones extracted by SigLIP. Once distilled, the semantic layers $\mathcal{S}(\cdot)$ are capable of extracting rich semantic features from both clean and noised visual latents. In statistics, the extracted features from clean visual latents by $\mathcal{S}(\cdot)$ have converged to an average cosine similarity of around 0.9 with those extracted by the original SigLIP on the curated 66M image-text pairs. We interpolate the position embeddings in the bicubic mode when involving other image/video resolutions.
Our models build upon two LLM variants, i.e., Qwen2.5-1.5B-Instruct [122] and Qwen2.5-7B-Instruct [122], respectively. We adopt 3D causal VAE proposed in Wan2.1 [100] with 8$\times$ and 4$\times$ spatial and temporal compression, respectively. In stage 1, we first train the 1.5B variant for 150K iterations using AdamW optimizer with a constant learning rate of 0.0001 on the curated 66M image-text pairs in a resolution of $432\times 432$. The context length of single image-text pairs is set as 1024. The total batch sizes for multimodal understanding and generation are 128 and 384, respectively. $\alpha$ in Eq. 4 is set as 0.2. For visual generation data, the caption is dropped with a probability of 0.1 to enable the classifier-free guidance. This training process roughly takes one and a half days using 64 H100 GPUs. Subsequently, we replace the generation data with 16M high-quality data (filtered from 66M image-text pairs) and continue to train for 40K iterations. In stage 2, we follow the training strategies in LLaVA-OneVision[52] to train the 1.5B model using around 9M multimodal instructional and 16M high-quality generation data for a total of around 35K iterations. $\alpha$ in Eq. 4 is set as 1.0. The stage 2 training process takes around 15 hours. For models with mixed-modality and video generation capabilities, we progressively add video-text and interleaved data in stage 1. For video data, we randomly sample a 2s 480p or 432$\times$432 clips with 17 frames from each video with an interval of 3 frames. The context length at this time is set as 7006. In stage 2, high-quality video-text and interleaved data are added to further improve video and mixed-modality generation capabilities.
@@
Table 3: Evaluation on multimodal understanding benchmarks. # Params. indicates the number of parameters of base LLM. * indicates the method uses two distinct models or sets of parameters for multimodal understanding and generation, respectively. Und. indicates “understanding”. Results in gray indicate the performance of und. only models or models with total parameters more than 13B.
@@
4.2 Multimodal Understanding
Quantitative Results. Table 3 highlights the performance of our models on multimodal understanding benchmarks, evaluated across metrics such as MME [37], GQA [46], SEED-Bench [51], MM-Bench [70], MMU [126], MMStar [20], and AI2D [48]. As shown in the table, both the 1.5B and 7B variants of our model consistently outperform state-of-the-art models across many metrics. For models with similar parameter sizes (1.5B), our model achieves the best scores on MME-p and MMU-val benchmarks while delivering competitive performance on GQA and SEED-Bench metrics. When compared to larger models with approximately 7B parameters, our models surpass state-of-the-art models such as Janus-Pro and even the significantly larger TokenFlow-XL model (14B parameters) in metrics including MME-p, GQA, MMMU-val, MMStar, and AI2D, while maintaining competitive performance on SEED-Bench and MM-Bench. These results underscore the robust perception capabilities of our unified visual representations, demonstrating their effectiveness in multimodal understanding tasks and the promising potentials in this domain.
@@ 4.3 Visual Generation
Image Generation. We compare our model with the state-of-the-art approaches on GenEval [39] and DPG-Bench [42] benchmarks in Tables 4 and 5. One can observe that our model surpasses most approaches, including TokenFlow-XL, Show-o, Emu3, and Transfusion, on the GenEval benchmark. Compared to Janus-Pro, which was trained on a significantly larger dataset of 144M image-text pairs,
@@
our model achieves promising results with only 66M image-text pairs. On DPG-Bench evaluation, our model has demonstrated the best overall score compared to generation-only models such as SD3-Medium and unified models, including Emu3-DPO and Janus-Pro. We also show qualitative results in Fig. 2 to illustrate that our model can generate high-quality and realistic images.
Video Generation. We compare our model with the text-to-video and image-to-video generation models in Tables 6 and 7. One can observe that with only 2B parameters, our model outperforms models such as Show-1, Emu3, and VILA-U with more than 6B parameters. Besides, our model has demonstrated competitive performance compared to CogVideoX and Step-Video-T2V. We also provide qualitative results of the text-to-video and image-to-video generation capability of our model in the middle of Fig. 2. One can observe that, given text prompts or an input image, our model can generate consistent video frames with reasonable motions, such as the smiling girl, lapping waves, and floating clouds.
@@ 4.4 Mixed-Modality Generation
Models # Params. Total QS SS SC BC TF MS DD AQ IQ OC MO HA C SR S AS TS OC’
@@
Table 6: Comparison with text-to-video models on the VBench [45] benchmark. # Params. indicates the number of total parameters for video generation including the base LLM and flow head. QS: Quality Score, SS: Semantic Score, SC: Subject Consistency, BC: Background Consistency, TF: Temporal Flickering, MS: Motion Smoothness, DD: Dynamic Degree, AQ: Aesthetic Quality, IQ: Imaging Quality, OC: Object Class, MO: Multiple Objects, HA: Human Action, C: Color, SR: Spatial Relationship, S: Scene, AS: Appearance style, TS: Temporal Style, OC’: Overall Consistency.
Models I2V Subject I2V Background Camera Motion Subject Consistency Background Consistency Temporal Flickering Motion Smoothness Dynamic Degree Aesthetic Quality Imaging Quality
@@ 4.5 Ablation Studies
Table 8: Effect of spatial (-temporal) fusion.
[…]
@@ 6 Conclusion
We appreciate Haozhe Liu for his consistent input and discussions regarding this project.
@@ Appendix A Technical Appendices and Supplementary Material
[…]
@@
[…]
@@
[…]
@@ A.2 Text Prompts
Table 12: Evaluation on the WISE [76] benchmark. # Params. indicates the number of parameters of the base LLM.
[…]
Native Unified Orthus-Instruct [50] 7B 0.23 0.31 0.38 0.28 0.31 0.20 0.27
Show-o [118] 1.3B 0.28 0.36 0.40 0.23 0.33 0.22 0.30
VILA-U [114] 7B 0.26 0.33 0.37 0.35 0.39 0.23 0.31
Janus-Pro [25] 1.5B 0.20 0.28 0.45 0.24 0.32 0.16 0.26
Janus-Pro [25] 7B 0.30 0.37 0.49 0.36 0.42 0.26 0.35
Show-o2 (Ours) 1.5B 0.30 0.38 0.47 0.32 0.42 0.28 0.35
Show-o2 (Ours) 7B 0.33 0.42 0.53 0.39 0.45 0.26 0.39
A.3 Additional Evaluation Results
We provide evaluation results on the WISE [76] benchmark in Table 12.

### v3

@@ 1 Introduction
Janus-Series [26, 79, 27]
AR (+Diff)
UnidFluid [38]
AR + MAR
Mogao [65]
AR + Diff.
@@
SEED-X [40]
AR + Diff.
ILLUME [111]
AR + Diff.
MetaMorph [106]
AR + Diff.
MetaQueries [83]
AR + Diff.
TokenFlow [89]
AR
@@ 3.1 Overall Framework
$\mathcal{L}_{\text{distill}}=-\frac{1}{n}\sum\log\text{sim}(\mathcal{S}(\textbf{x}_{t}),\text{\tt{SigLIP}}(\textbf{X})),$ (2)
@@ 3.2 Training Recipe
Stage-1 Projector 66M
WebVid [8]
Pandas [23]
OmniCorpus [60]
@@
Stage-2 Full Model (w/o VAE) 9M HQ Und. OpenVid-1M [80] Gen. VIST [47]
16M HQ Gen. 1.5M Internal Data Gen. CoMM [24]
1.6M Video Und.
@@
Stage-2. Subsequently, we tune the full model using 9M high-quality multimodal understanding instruction data, 16M high-quality visual generation data filtered from the 66M image-text pairs, and 1.6M video understanding data.
@@ 4.1 Experimental Setup
Datasets. The curated approximately 66M image-text pairs consist of images with a resolution of at least 512 pixels in width and height. The images are filtered from CC12M [14], COYO [12], LAION-Aesthetic-12M∗* ∗* * https://huggingface.co/datasets/dclure/laion-aesthetics-12m-umap and AI synthetic data. The images are recaptioned by LMMs except for the synthetic data. The 9M high-quality multimodal understanding instruction data is curated from Densefusion-1M [62], and LLaVA-OneVision [56].
Implementation Details. The semantic layers $\mathcal{S}(\cdot)$ are pre-distilled from SigLIP-so400m-patch14-384∗* ∗* * https://huggingface.co/google/siglip-so400m-patch14-384 over 200K iterations, using a batch size of 512 and a cosine-scheduled learning rate of 2e-5. During distillation, Eq. 1 is applied to the visual latents with only a probability of 0.3 in the last 20K iterations. The input image resolution of 3D causal VAE encoder with $2\times 2$ patch embedding layer is set as $432\times 432$ to get $729=27\times 27$ visual latents, which matches the ones extracted by SigLIP. Once distilled, the semantic layers $\mathcal{S}(\cdot)$ are capable of extracting rich semantic features from both clean and noised visual latents. In statistics, the extracted features from clean visual latents by $\mathcal{S}(\cdot)$ have converged to an average cosine similarity of around 0.9 with those extracted by the original SigLIP on the curated 66M image-text pairs. We interpolate the position embeddings in the bicubic mode when involving other image/video resolutions.
[…]
@@
Building on the pre-trained image-level Show-o2 models, we enhance their video understanding capabilities by further fine-tuning on 1.6M video samples from [145], together with 1.1M image-level samples from the earlier stage. We adopt the same video training and inference settings as LLaVA-OneVision. The evaluation results are shown in Table 4.
Table 3: Evaluation on multimodal understanding benchmarks. # Params. indicates the number of parameters of base LLM. * indicates the method uses two distinct models or sets of parameters for multimodal understanding and generation, respectively. indicates the Show-o2 models fine-tuned using video understanding data. Und. indicates “understanding”. Results in gray indicate the performance of und. only models or models with total parameters more than 13B.
@@
4.2 Multimodal Understanding on Images and Videos
Quantitative Results. Table 3 highlights the performance of our models on multimodal understanding benchmarks, evaluated across metrics such as MME [39], GQA [49], SEED-Bench [55], MM-Bench [75], MMU [137], MMStar [21], and AI2D [51]. As shown in the table, both the 1.5B and 7B variants of our model consistently outperform state-of-the-art models across many metrics. For models with similar parameter sizes (1.5B), our model achieves the best scores on MME-p and MMU-val benchmarks while delivering competitive performance on GQA and SEED-Bench metrics. When compared to larger models with approximately 7B parameters, our models surpass state-of-the-art models such as Janus-Pro and even the significantly larger TokenFlow-XL model (14B parameters) in metrics including MME-p, GQA, MMMU-val, MMStar, and AI2D, while maintaining competitive performance on SEED-Bench and MM-Bench. These results underscore the robust perception capabilities of our unified visual representations, demonstrating their effectiveness in multimodal understanding tasks and the promising potential in this domain. In addition, we present the video understanding performance of Show-o2 in Table 4.
Table 4: Evaluation on video understanding benchmarks. # Params. denotes the number of parameters in the base LLM, while # Frames represents the maximum number of video frames used during training and inference. Und. stands for understanding. marks the Show-o2 models that have been fine-tuned on video understanding data. All results are reported in terms of zero-shot accuracy.
Model # Params. # Frames
ActNet-QA
MVBench
NExT-QA
PerceptionTest
LongVideoBench
VideoMME
test test mc val val wo/w-subs
Proprietary Und. Only Models
GPT-4V [81] - - 57.0 43.5 - - 61.3 59.9/63.3
GPT-4o [82] - - - - - - 66.7 71.9/77.2
Gemini-1.5-Flash [103] - - 55.3 - - - 61.6 70.3/75.0
Gemini-1.5-Pro [103] - - 57.5 - - - 64.0 75.0/81.3
Open-source Und. Only Models
VILA [69] 40B - 58.0 - 67.9 54.0 - 60.1/61.1
PLLaVA [131] 34B 16 / 16 60.9 58.1 - - 53.2 -
LongVA [143] 7B - 50.0 - 68.3 - - 52.6/54.3
IXC-2.5 [142] 7B 64 / 64 52.8 69.1 71.0 34.4 - 55.8/58.8
LLaVA-OV [56] 7B 32 / 32 56.6 56.7 79.4 57.1 56.5 58.2/61.5
VideoLLaMA2 [30] 7B 16 / 16 50.2 54.6 - 51.4 - 47.9/50.3
Unified Multimodal Models
Show-o2(Ours) 1.5B 32 / 32 52.7 49.8 72.1 56.1 49.2 48.0/51.6
Show-o2(Ours) 7B 16 / 32 56.4 55.8 79.0 61.9 55.5 57.4/60.9
@@ 4.2 Multimodal Understanding on Images and Videos
Table 7: Overall quantitative comparison of different methods on OneIG-Bench. Gen. denotes “generation”. # Params. indicates the number of parameters of base LLM. # Data. indicates the number of image-text pairs used for visual generation during training.
[…]
Gen. Only SD3.5-Large[37] 8B - 0.809 0.629 0.294 0.353 0.225
Flux.1-dev [54] 12B - 0.786 0.523 0.253 0.368 0.238
SANA-1.5 (PAG) [126] 4.8B - 0.765 0.069 0.217 0.401 0.216
Lumina-Image 2.0 [88] 2.6B 110M 0.819 0.106 0.270 0.354 0.216
HiDream-I1-Full [44] 17B - 0.829 0.707 0.317 0.347 0.186
Unified Models Show-o-512 [128] 1.3B 2B 0.702 0.002 0.213 0.361 0.241
Janus-Pro [27] 7B 144M 0.553 0.001 0.139 0.276 0.365
BLIP3-o [18] 8B 55M 0.711 0.013 0.223 0.361 0.229
BAGEL [32] 14B 1600M 0.769 0.244 0.173 0.367 0.251
OmniGen2 [117] 7B 150M 0.804 0.680 0.271 0.377 0.242
Show-o2 (Ours) 1.5B 66M 0.798 0.002 0.219 0.317 0.186
Show-o2-$\times1024$ (Ours) 1.5B 66M 0.798 0.125 0.274 0.351 0.186
Show-o2 (Ours) 7B 66M 0.817 0.002 0.226 0.317 0.177
@@ 4.3 Visual Generation
Image Generation. We compare our model with the state-of-the-art approaches on GenEval [41], DPG-Bench [45], and OneIG[13] benchmarks in Tables 5, 6, and 7. One can observe that our model surpasses most approaches, including TokenFlow-XL, Show-o, Emu3, and Transfusion, on the GenEval benchmark. Compared to Janus-Pro, which was trained on a significantly larger dataset of 144M image-text pairs, our model achieves promising results with only 66M image-text pairs. On DPG-Bench evaluation, our model has demonstrated the best overall score compared to generation-only models such as SD3-Medium and unified models, including Emu3-DPO and Janus-Pro. On OneIG-Bench, our models also achieve competitive performance. We also show qualitative results in Fig. 2 to illustrate that our model can generate high-quality and realistic images.
Video Generation. We compare our model with the text-to-video and image-to-video generation models in Tables 8 and 9. One can observe that with only 2B parameters, our model outperforms models such as Show-1, Emu3, and VILA-U with more than 6B parameters. Besides, our model has demonstrated competitive performance compared to CogVideoX and Step-Video-T2V. We also provide qualitative results of the text-to-video and image-to-video generation capability of our model
@@
in the middle of Fig. 2. One can observe that, given text prompts or an input image, our model can generate consistent video frames with reasonable motions, such as the smiling girl, lapping waves, and floating clouds.
@@ 4.4 Mixed-Modality Generation
Table 8: Comparison with text-to-video models on the VBench [48] benchmark. # Params. indicates the number of total parameters for video generation including the base LLM and flow head. QS: Quality Score, $S$: Semantic Score, SC: Subject Consistency, BC: Background Consistency, TF: Temporal Flickering, MS: Motion Smoothness, $D$: Dynamic Degree, AQ: Aesthetic Quality, IQ: Imaging Quality, OC: Object Class, MO: Multiple Objects, HA: Human Action, C: Color, SR: Spatial Relationship, S: Scene, AS: Appearance style, TS: Temporal Style, OC’: Overall Consistency.
Models # Params. Total QS $S$ SC BC TF MS $D$ AQ IQ OC MO HA C SR S AS TS OC’
@@
Models
I2V
Subject
I2V
Background
Camera
Motion
Subject
Consistency
Background
Consistency
Temporal
Flickering
Motion
Smoothness
Dynamic
Degree
Aesthetic
Quality
Imaging
Quality
@@ 4.5 Ablation Studies
Table 10: Impact of spatial (-temporal) fusion.
MMEp-$\text${p} $\uparrow$ GQA $\uparrow$ POPE $\uparrow$ FID-5K $\downarrow$
@@
Table 13: Impact of training recipe on text-only performance. One-stage training denotes full-parameter-co-training on image-text pairs and the text-only RefinedWeb [87] data. Note that the curated multimodal understanding data consists of text-only instructional data. We perform the evaluation under the same setting using the lm-evaluation-harness tool.
Models # Params. Training Recipe MMLU GPQA GSM8K HumanEval
Qwen2.5 Instruct [133] 1.5B - 60.20 $\pm$ 0.39 28.12 $\pm$ 2.13 51.86 $\pm$ 1.38 35.37 $\pm$ 3.74
Show-o2 (Ours) 1.5B One-stage training with RefinedWeb 28.25 $\pm$ 0.38 25.00 $\pm$ 2.05 4.55 $\pm$ 0.57 3.05 $\pm$ 1.35
Show-o2 (Ours) 1.5B Our two-stage training 56.75 $\pm$ 1.37 29.24 $\pm$ 2.15 49.43 $\pm$ 1.38 35.54 $\pm$ 3.70
Qwen2.5 Instruct [133] 7B - 71.75 $\pm$ 0.36 32.37 $\pm$ 2.21 82.49 $\pm$ 1.05 65.24 $\pm$ 3.73
Show-o2 (Ours) 7B One-stage training with RefinedWeb 28.43 $\pm$ 0.21 26.34 $\pm$ 2.08 1.52 $\pm$ 0.34 4.01 $\pm$ 1.25
Show-o2 (Ours) 7B Our two-stage training 70.73 $\pm$ 0.36 31.47 $\pm$ 2.22 75.28 $\pm$ 1.19 70.73 $\pm$ 3.56
Table 13 shows that our models effectively preserve language knowledge and achieve performance comparable to the original Qwen2.5-1.5B and Qwen2.5-7B Instruct models. In contrast, direct one-stage full-parameter-co-training with textual data such as RefinedWeb results in substantial performance degradation, highlighting the necessity of the two-stage training approach when high-quality corpora are unavailable.
Table 14: Impact of image token count on chart, text, and document VQA.
[…]
LLaVA-OV 7B 729 56.24 62.71 39.59 66.19
Show-o2 7B 729 48.00 59.34 42.31 62.92
[…]
As shown in Table 14, our ablation study reveals that increasing the number of image tokens significantly boosts performance across all tasks, even though the model was trained with a fixed image resolution. Using the AnyRes strategy at inference time consistently improves results, highlighting the benefit of higher token counts for capturing fine-grained details. When compared to the baseline LLaVA-OV-7B, our model achieves comparable results on DocVQA, InfoVQA, and TextVQA validation sets, but underperforms on ChartQA. We attribute this gap to the limited chart-related data available during semantic layer distillation, which constrains the model’s ability to capture chart-specific information. We believe that incorporating more OCR and document-centric data into the distillation process will further strengthen the unified model’s OCR and document understanding capabilities.
@@ 6 Conclusion
We thank Haozhe Liu for his valuable input and discussions throughout this project. We are also grateful to Meng Wei and Weihao Wang for their assistance in preparing and organizing the datasets for image and video generation.
@@ Appendix A Technical Appendices and Supplementary Material
MonoFormer [146]
AR + Diff.
Dual-Diffusion [63]
Diff.
SynerGen-VL [58]
AR
MMAR [134]
AR + MAR
MUSE-VL [129]
AR
Orthus [53]
AR + Diff.
Liquid [118]
AR
LlamaFusion [95]
AR + Diff.
UGen [99]
AR
UniDisc [98]
Diff.
UniToken [50]
AR
Harmon [122]
AR+MAR
DualToken [96]
AR
UniTok [77]
AR
Selftok [110]
AR
Muddit [94]
Diff.
MMaDA [135]
Diff.
@@
Janus-Series [26, 79, 27]
AR (+Diff.)
VARGPT [148]
AR
UnidFluid [38]
AR + MAR
OmniMamba [149]
AR
Mogao [65]
AR + Diff.
@@
MetaMorph [106]
AR + Diff.
ILLUME [111]
AR + Diff.
ILLUME+ [46]
AR + Diff.
MetaQueries [83]
AR + Diff.
Nexus-Gen [141]
AR + Diff.
Ming-Lite-Uni [42]
AR + Diff.
BLIP3-o [18]
AR + Diff.
OpenUni [121]
AR + Diff.
UniWorld [67]
AR + Diff.